# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v1

@@ VI-D Takeaways
In this section, we categorize two major agent-agent communication protocols architectures: CS-based and P2P-based. Accordingly, the security risks are also multifaceted: CS-based architecture put heavy burdens on the centralized agent servers, such as registration pollution and SEO poisoning. P2P-based architecture suffer from the lack of efficient and centralized management of agents, such as non-convergence and man-in-the-middle attacks. Besides, both of them are vulnerable to universal risks, such as agent spoofing, bullying, and privacy leakage. We also discussed potential defenses countermeasures targeting each risk. We believe that as agent-agent communication continues to grow, more vulnerabilities in this process will be discovered..
@@ VII Agent-Environment Communication
This section begins by reviewing key protocol designs that enable compositional and standardized communication between agents and environments, then examines the associated security risks, including vulnerabilities in memory, retrieval-augmented reasoning, tool invocation, and multi-tool workflows. Finally, We outlook on the defense strategies for mitigating these threats and securing agent-environment interaction.
@@ VII-A1 MCP
MCP adopts a modular architecture comprising three core components: the host, the client, and the server. The host functions as a trusted local orchestrator responsible for managing the lifecycle of clients, enforcing access control policies, and mediating secure interactions in potentially multi-tenant environments. The client represents the interaction thread of a specific agent or session. It discovers available tools, formulates structured invocations, and handles synchronous or asynchronous responses during task execution. The server serves as a centralized registry that maintains and exposes tool specifications, contextual prompts, and workflow templates. These tools can follow either a declarative pattern (e.g., describing operations such as information retrieval) or an imperative pattern (executing executable calls like SQL queries or document edits).
@@ VII-A3 Function Calling Mechanisms
VII-B Security Risk Analysis
@@
As the capabilities of LLM-powered agents continue to evolve, their interactions with the external world become increasingly complex and powerful. In particular, the integration of memory systems and external tool invocation-two key enablers of persistent, autonomous behavior-introduce a new set of attack surfaces that adversaries can exploit. This section provides an in-depth analysis of the security risks that arise specifically from these two modules: the memory module, responsible for storing and retrieving contextual information, and the tool module, which enables agents to execute actions by interfacing with external systems or services (e.g., via function calls). We first explain how these two components typically function in LLM-agent ecosystems, and outline the general attack paradigms against each. We then provide detailed analyses of specific vulnerabilities, attack techniques, and representative works from the security literature that highlight these threats.
VII-B1 Memory-based and RAG-based Risks
@@
Memory-based Risks. Memory modules play a crucial role in enabling LLM-based agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [332]. Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term information through external storage systems, such as vector databases or document repositories. These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [332].
@@
Typically, a memory module operates through three stages: write, retrieve, and apply. During the write phase, the agent logs past utterances, tool outputs, subgoals, or retrieved facts into memory. Later interactions initiate the retrieve phase, where semantically similar records are fetched via embedding matching or keyword search. These records are then injected into the model’s context window or used for downstream decisions, forming the apply phase. While this architecture empowers agents with dynamic reasoning abilities, it also introduces new vulnerabilities that extend beyond the conventional LLM prompt space.
@@
Memory Injection. In memory injection attacks, adversaries insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access. The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content. Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations. A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [57, 288]. These steps, once embedded in memory, become semantically linked to a targeted victim query. When the victim issues a benign instruction, the poisoned memory is likely to be retrieved, thereby hijacking the agent’s planning process. This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
@@
Memory Poisoning. Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads. These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered. During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs. Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [34]. This method generalizes across agent types and does not require model access or parameter modification.
@@
Memory Extraction. In addition to injection and poisoning, memory modules pose risks of unintended information leakage. Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data. This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [313, 118]. It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [260]. Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces. These findings highlight not only the retrieval vulnerability, but also the insufficiency of downstream response filtering as a defense.
Real-world Consequences. The practical implications of compromised memory are non-trivial, especially when coupled with autonomous execution capabilities. For example, targeted poisoning of domain-specific memory has led to agents generating toxic chemical synthesis plans under the influence of forged scientific references [152]. In such cases, malicious records retrieved during planning phases corrupted the reasoning chain of the agent, triggering hazardous tool calls. These examples underscore the dangerous entanglement between corrupted memory and downstream actions, particularly in scientific, medical, or high-stakes decision-making domains.
@@
Retrieval-Augmented Generation (RAG) combines the generative strength of large language models (LLMs) with the factual accuracy and relevance of external information retrieval systems. Instead of relying solely on parametric knowledge stored within the pretrained model, RAG augments generation by sourcing passages from an external knowledge base in response to the input query. These retrieved documents are then concatenated with the query and passed into the LLM for final answer generation. This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
@@
Knowledge Corruption via Data Poisoning. A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries. These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content. Once injected into the knowledge base, they can be prioritized during retrieval and directly influence the LLM’s final response.
Several recent works have demonstrated the feasibility of such attacks. PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [348]. Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [197]. Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [341].
Privacy Risks and Unintended Leakage. RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs. This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus. The risk is amplified when access permissions on the corpus are loosely controlled or aligned purely through similarity metrics.
Recent studies have called attention to this concern. Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [313]. These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
Broader Threat Surface and Evaluation Gaps. Compared to memory modules, RAG corpora are often larger, dynamically updatable, and more difficult to monitor. Because retrieval corpora may be sourced from web documents, community-shared datasets, or user uploads, attackers can often poison them without interacting directly with the agent. Moreover, dense retrieval introduces additional attack vectors via embedding collisions or adversarial representation alignment, wherein malicious documents are optimized to collide with benign queries in the retriever’s latent space.
VII-B2 Tool-based Risks
@@
Malicious Tools as Attack Vectors Given that many tools are authored externally or retrieved from shared tool repositories, attackers may publish seemingly benign tools containing covert malicious logic. Beyond executable payloads, adversaries often embed hidden prompts or jailbreaking instructions in tool metadata fields such as descriptions, example usages, or API annotations. These embedded messages can influence the LLM’s planning behavior, bypass output constraints, or redirect queries.
For instance, malicious tools may be structured to leak user inputs, intercept queries, or bias tool choices. Prior work has demonstrated that embedding adversarial cues into tool schemas enables agents to select harmful tools under non-malicious queries [6]. Such “trojan tooling” exploits the model’s trust in structured function interfaces and the absence of output sanitization or behavioral filters in many open frameworks.
Misuse of Legitimate Tools Even when tools themselves are benign, adversaries may manipulate the agent into misusing them through crafted inputs, prompt injections, or indirect instructions. Common threats include API misuse, argument-level command injection, and unsafe rendering behaviors. For example, LLM agents with Markdown preview tools may inadvertently expose internal strings such as emails or credentials by embedding malicious image links within generated output [71]. These links, when auto-rendered by browsers or viewers, initiate unintended HTTP requests that leak private context to attacker-controlled servers.
More generally, tools that perform unsafe network, file, or system actions can be exploited using crafted arguments [306]. Improper validation may lead to server-side request forgery (SSRF) [18], arbitrary file access, code injection, or data exfiltration. In multi-tenant use cases, a successful tool misuse may compromise session boundaries or extract environment-level secrets such as API keys, tokens, or runtime credentials.
@@
Manipulation of the Tool Selection Process Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation. This selection logic can be hijacked. Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
@@
Research indicates that adversaries can generate synthetic tool descriptions that stealthily override the model’s planning process [237]. These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations. Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during planning phase. Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
Cross-Tool Chaining Exploits As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls. These workflows-e.g., summarize(search(”…”))-blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations. Without inter-tool validation, an adversary can exploit pipeline dependencies to propagate malicious content downstream.
@@
Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools). In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server,” which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [40].
Such “retrieval-agent deception” demonstrates a broader class of attacks where loosely regulated multi-tool interactions allow adversarial instructions to percolate through tool output and trigger unintended behaviors. When tools are pipelined programmatically-without adequate sanitization or output auditing-malicious payloads can cross logical layers and become effective system-level exploits.
VII-C Defense Countermeasure Prospect

### v2

@@ VI-D Takeaways
In this section, we categorize two major agent-agent communication architectures: CS-based and P2P-based. Accordingly, the security risks are also multifaceted: CS-based architecture puts heavy burdens on the centralized agent servers, such as registration pollution and SEO poisoning. P2P-based architecture suffers from the lack of efficient and centralized management of agents, such as non-convergence and man-in-the-middle attacks. Besides, both of them are vulnerable to universal risks, such as agent spoofing, bullying, and privacy leakage. We also discussed potential defense countermeasures targeting each risk. The defenses discussed in this section can also mitigate the risks from compromised agents against user (Section V-C) and environments (Section VII-C). We believe that as agent-agent communication continues to grow, more vulnerabilities in this process will be discovered.
@@ VII Agent-Environment Communication
This section begins by reviewing key protocol designs that enable compositional and standardized communication between agents and environments, then examines the associated security risks, including vulnerabilities in memory, retrieval-augmented reasoning, tool invocation, and multi-tool workflows. Finally, we discuss the defense strategies for mitigating these threats and securing agent-environment communication.
@@ VII-A1 MCP
As shown in Figure 9, MCP adopts a modular architecture comprising three core components: the host, the client, and the server. The host functions as a trusted local orchestrator responsible for managing the lifecycle of clients, enforcing access control policies, and mediating secure interactions in potentially multi-tenant environments. The client represents the interaction thread of a specific agent or session. It discovers available tools, formulates structured invocations, and handles synchronous or asynchronous responses during task execution. The server serves as a centralized registry that maintains and exposes tool specifications, contextual prompts, and workflow templates. These tools can follow either a declarative pattern (e.g., describing operations such as information retrieval) or an imperative pattern (executing executable calls like SQL queries or document edits).
@@ VII-A3 Function Calling Mechanisms
TABLE VII: The Security Risks in the Agent-Environment Communication Phase and Their Characteristics
Stage Risk Type Representative Threats Attack Characteristics
Malicious environments against benign agents Memory-related Risks Memory Injection Induce agents to generate and record harmful content through natural interactions
Memory Poisoning Implant pairs of examples containing adversarial triggers and payloads
Memory Extraction Extract sensitive data in memory modules without explicit queries for privacy
Knowledge-related Risks Knowledge Corruption Inject adversarial texts into external knowledge bases to influence the response
Privacy Risks Craft prompts to induce agents to recover sensitive or private content
Tool-related Risks Malicious Tools Malicious tool functions or prompts embedded in the tool description
Manipulation of Tool Selection Inject misleading prompts to make agents select the compromised tool
Cross-Tool Chaining Exploits Manipulate the output of a tool to inject malicious inputs for subsequent tools
Compromised agents against benign environments Information Corruption Propagate Tampered Information Undermine an agent’s memory and knowledge using misleading information
Misusing Tools Data Exfiltration Extract and transfer confidential information out of protected environments
System and Service Disruption Disrupt the functionality of services without directly targeting data confidentiality
Malicious Content Propagation Spread malicious content to trusted storage resources
Real-World Damage Digital Environment Contamination Introduce harmful elements into digital ecosystems, thereby corrupting integrity
Physical Environment Disruption Compromised agents cause disruptions to real-world physical environments
VII-B Security Risk Analysis: Malicious Environments Against Benign Agents
@@
As the capabilities of LLM-powered agents continue to evolve, their interactions with the external world become increasingly complex and powerful. In particular, the integration of memory systems and external tool invocation introduces a new set of attack surfaces that attackers can exploit. This section provides an in-depth analysis of the security risks that arise specifically from these two modules: the memory module, responsible for storing and retrieving contextual information, and the tool module, which enables agents to execute actions by interfacing with external systems or services (e.g., via function calls). We first explain how these two components typically function in the agent ecosystem and outline the general attack paradigms against each. We then provide detailed analyses of specific vulnerabilities, attack techniques, and representative works from the security literature that highlight these threats.
VII-B1 Memory-related Risks
@@
Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [351]. Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term historical information through external storage systems, such as vector databases or document repositories. These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [351].
@@
Typically, a memory module operates through three stages: write, retrieve, and apply. During the write phase, the agent logs past utterances, tool outputs, subgoals, or retrieved facts into memory. Later interactions initiate the retrieval phase, where semantically similar records are fetched via embedding matching or keyword search. These records are then injected into the model’s context window or used for downstream decisions, forming the apply phase. While this architecture empowers agents with dynamic reasoning abilities, it also introduces new vulnerabilities that extend beyond the conventional LLM prompt space.
@@
  • Memory Injection: In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access. The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content. Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations. A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [61, 307]. These steps, once embedded in memory, become semantically linked to a targeted victim query. When the victim issues a benign instruction, the poisoned memory is likely to be retrieved, thereby hijacking the agent’s planning process. This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
@@
  • Memory Poisoning: Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads. These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered. During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs. Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [38]. This method generalizes across agent types and does not require model access or parameter modification.
@@
  • Memory Extraction: In addition to injection and poisoning, memory modules pose risks of unintended information leakage. Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data. This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [332, 128]. It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [278]. Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces. These findings highlight not only the retrieval vulnerability but also the insufficiency of downstream response filtering as a defense.
VII-B2 Knowledge-related Risks
@@
External knowledge techniques, such as Retrieval-Augmented Generation (RAG), combine the generative strength of LLMs with the factual accuracy and relevance of external knowledge retrieval systems. Instead of relying solely on parametric knowledge stored within the pretrained model, RAG augments generation by sourcing passages from an external knowledge base in response to the input query. These retrieved documents are then concatenated with the query and passed into the LLM for final answer generation. This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
@@
  • Knowledge Corruption via Data Poisoning: A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries. These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content. Once injected into the knowledge base, they can be prioritized during retrieval and directly influence the LLM’s final response. Several recent works have demonstrated the feasibility of such attacks. PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [365]. Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [209]. Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [358].
  • Privacy Risks and Unintended Leakage: RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs. This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus. The risk is amplified when access permissions on the corpus are loosely controlled or aligned purely through similarity metrics. Recent studies have called attention to this concern. Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [332]. These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
Compared to memory modules, RAG corpora are often larger, dynamically updatable, and more difficult to monitor. Because retrieval corpora may be sourced from web documents, community-shared datasets, or user uploads, attackers can often poison them without interacting directly with the agent. Moreover, dense retrieval introduces additional attack vectors via embedding collisions or adversarial representation alignment, wherein malicious documents are optimized to collide with benign queries in the retriever’s latent space.
VII-B3 Tool-related Risks
@@
  • Malicious Tools as Attack Vectors: Given that many tools are authored externally or retrieved from shared tool repositories, attackers may publish seemingly benign tools containing covert malicious logic. For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [111] but also tool metadata fields such as descriptions, example usages, or API annotations[254, 72, 101, 150, 229]. These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries.
@@
  • Manipulation of Tool Selection: Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation. This selection logic can be hijacked. Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
@@
    Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [250]. These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations. Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase. Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
@@
  • Cross-Tool Chaining Exploits: As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls. These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations. Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools). In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [44].
VII-C Security Risk Analysis: Compromised Agents Against Benign Environment
In Section VII-B, we analyzed how the malicious environments can influence an agent through its memory and tool modules. However, once an agent is compromised, it can cause widespread and lasting damage to external environments. This section explores this reverse threat, focusing on how compromised agents can exploit their functionalities to contaminate memory modules, abuse external tools, and pose risks to the real world.
VII-C1 Corrupting Memory and Knowledge
Memory and knowledge bases are central to how the agent thinks and acts. Once these components are corrupted, they can lead to incorrect reasoning and decisions. Moreover, the damage can infect other agents and cause long-term harm to the entire system.
A compromised agent can become a source of systemic contamination. Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [133, 38, 158]. Once the shared knowledge base is compromised, other agents may unknowingly retrieve and integrate malicious information into their memory module during task execution, converting knowledge base contamination into memory infection. Subsequently, an agent with a compromised memory module can use its authorized write access to contaminate the shared knowledge base of the entire system, in turn, forming a reverse contamination loop from memory to knowledge. Since these contamination operations come from the trusted agents within the system, they are highly difficult to detect. Once established, the contamination can persist long-term, continuously disrupting the behavior of compromised agents and misleading other agents that rely on the same knowledge sources, resulting in a form of chronic poisoning of the information ecosystem.
VII-C2 Misusing Tools
Tools serve as the channels through which an agent interacts with the world. Once the agent is compromised, tools can be exploited to cause harm. Since agents are often granted permissions to execute specific API calls or system commands, misuse of these tools can lead to severe consequences.
  • Data Exfiltration: If a compromised agent has tool permissions to access databases, file systems, or communication interfaces, it can be manipulated to extract and transmit confidential data, intellectual property, or personally identifiable information. These payloads can be covertly delivered to external servers via email, HTTP requests, or chained tool invocations. For instance, a compromised agent equipped with Markdown rendering capabilities might unknowingly embed malicious image links in generated content. When rendered by a browser, these links trigger covert HTTP requests that exfiltrate sensitive data such as email addresses or access tokens to attacker-controlled servers. Notably, attackers may also exploit tool misuse to gain indirect access to backend systems and gradually escalate privileges through cross-toolchain interactions.
  • System and Service Disruption: A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to carry out destructive operations. A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to execute destructive operations. Attackers may exploit these privileges to launch internal network scans, initiate denial-of-service attacks, or issue commands, potentially causing system crashes or irreversible business data loss. Such actions are especially dangerous when triggered via tool interfaces that lack proper input validation. Vulnerabilities commonly exploited in these contexts include parameter injection, server-side request forgery (SSRF) [22], and arbitrary file access. Beyond technical damage, the compromised agent can be manipulated to disrupt business operations, such as erroneously invoking the “cancel all orders” tool. Moreover, adversaries may use retrieval tools to fetch misleading information, then exploit decision-making tools such as trading agents to automatically execute harmful operations, resulting in major financial losses.
  • Malicious Content Propagation: Agents with permissions to publish externally (e.g., via email, social media, or CMS APIs) can be exploited to widely spread malware, phishing links, or misinformation once compromised. For example, a trusted customer service agent may send malware-laden emails to clients, or a content generation agent might publish misleading articles on official websites. Because these agents are trusted entities, such attacks are highly deceptive. More dangerously, attackers can leverage access to contacts, email histories, and user preferences to craft highly personalized phishing campaigns, enabling large-scale social engineering attacks.
VII-C3 Real-World Damage
Compromised agents can cause harm far beyond their internal systems, affecting both digital and physical ecosystems through their interactions with the external world. This section examines how such agents may disrupt the real world by polluting shared digital environments and executing harmful physical actions.
  • Digital Environment Contamination: A compromised agent can cause long-term damage to the external digital environment, not by directly attacking other agents, but by polluting the shared information ecosystem they rely on. Since agents often interact with external platforms (such as submitting code on GitHub or editing Wikipedia entries), once compromised, they can systematically inject subtle yet harmful errors or biases into these shared resources [173]. Unlike cross-agent contamination, digital environment contamination indirectly infects all agents dependent on the polluted information sources. For example, a compromised coding agent may embed hidden logic bugs or backdoors when contributing code; a corrupted knowledge management agent might alter Wikipedia pages or internal knowledge bases by falsifying citations or inserting biased descriptions, thereby damaging the entire knowledge graph with far-reaching consequences.
  • Physical Environment Disruption: Once an agent’s memory or tool module is compromised, the resulting threat is not limited to digital risk. It can manifest as concrete damage to the physical world through specific decision-making chains and execution paths. Polluted memory may contain falsified sensor data, misleading the agent’s perception of the physical environment. Tool modules that serve as interfaces to physical systems can directly implement flawed decisions, affecting device behaviors, environmental control, or industrial processes. For example, an agricultural agent misled by false pest-related memories may overapply pesticides; a quality inspection robot referencing corrupted standard images may repeatedly approve defective components; a warehouse robot using a compromised path-planning module may unknowingly cause stacking imbalances and logistics bottlenecks. Critically, such behaviors often appear to follow normal procedures, making them difficult to detect through traditional logging or anomaly detection methods. Thus, the consequences of agent compromise extend beyond digital misinformation, posing a tangible risk of physical system disruption.
VII-D Defense Countermeasure Prospect